Find the first three iterates of the function f(x)=x^2-2, if the initial value is -2
Thank you so much for all of your help!!!!!!!
God Bless!!!!!!!

Answer by stanbon(75887)

Find the first three iterates of the function
f(x)=x^2-2, if the initial value is -2.
----------
f(-2) = (-2)^2-2 = 4-2 = 2
---
f(2) = 2^2-2 = 2
---
f(2) = 2^2-2 = 2
==================
Cheers,
Stan H.